OpenStudy (anonymous):

1+sin2θ / 1-sin2θ = tan^2 ( π/4 + θ)

OpenStudy (anonymous):

So prove it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Here use; \[\sin(2 \theta) = 2\sin(\theta)\cos(\theta)\] \[1 = \sin^2 ( \theta) + \cos^2 (\theta)\]

OpenStudy (anonymous):

\[\frac{1 + \sin(2 \theta)}{1 - \sin(2 \theta)} = \frac{\sin^2 (\theta) + \cos^2 (\theta) + 2\sin(\theta)\cos(\theta)}{\sin^2 (\theta) + \cos^2 (\theta) - 2\sin(\theta)\cos(\theta)} \implies \frac{(\sin(\theta) + \cos(\theta))^2}{(\sin(\theta) - \cos(\theta))^2}\]

OpenStudy (anonymous):

\[(\frac{\sin(\theta) + \cos(\theta)}{\sin(\theta) - \cos(\theta)})^2\]

OpenStudy (anonymous):

okay .. then ?

OpenStudy (anonymous):

Now divide by cos(theta) in the numerator and denominator: \[(\frac{1 +\tan(\theta)}{1 - \tan(\theta)})^2 \implies (\frac{\tan(45) + \tan(\theta)}{1 - \tan(45) \tan(\theta)})^2 \implies \tan^2(45 + \theta)\]

OpenStudy (anonymous):

\[45^{\circ} = \frac{\pi}{4}\]

OpenStudy (anonymous):

\[\tan(45) = 1\]

OpenStudy (anonymous):

(1+tan(θ)1−tan(θ))^2 how did yu get this ?

OpenStudy (anonymous):

\[\large (\frac{\frac{\sin(\theta)}{\cos(\theta)} + \frac{\cos(\theta)}{\cos(\theta)}}{\frac{\sin(\theta)}{\cos(\theta)} - \frac{\cos(\theta)}{\cos(\theta)}})^2 \implies (\frac{\tan(\theta) + 1}{\tan(\theta) -1 })^2\]

OpenStudy (anonymous):

In square you can change the order: so you can reverse the denominator: \[(\frac{\tan(\theta) + 1}{1 - \tan(\theta) })^2\]

OpenStudy (anonymous):

okay thank yu soo mch .. wish i had more medals for yu :D

OpenStudy (anonymous):

It is about knowledge and not medals.. Welcome dear..

OpenStudy (anonymous):

:)